# Name That Part: 3D Part Segmentation and Naming

### v1

@@ 5. Limitations and Future Work
Future work will focus on mitigating this dependency by exploring self-supervised or weakly supervised formulations and by incorporating stronger 3D priors, for example from generative models or skeletal/medial representations. Another important direction is to reduce the current reliance on frozen PartField features by enabling full end-to-end training of the geometric feature extractor, which was not pursued here primarily due to computational constraints rather than methodological ones. Despite these limitations, our framework is immediately usable by organizations and research labs with abundant compute and proprietary 3D assets, who can scale ALIGN-Parts to richer, closed-source datasets and drive progress towards truly large-scale 3D scene understanding at the part level.
6. Discussion
@@
ALIGN-Parts reframes semantic 3D part segmentation as a set alignment problem, where Partlets trained via optimal transport jointly learn geometry and semantics without part count supervision. Unlike Find3D’s per-point alignments or PartField’s brittle clustering, our end-to-end matching produces fast, coherent, named parts directly. LLM-generated affordance descriptions ("door handle is grasped to open a door") are important for disambiguating fine-grained parts that confuse simple part names, and MPNet’s handling of long-form text outperforms CLIP/SigLIP. Despite using 10× fewer points than PartField, we achieve superior fine-part localization (e.g., a scissors screw), confirming that semantic part-level representations are more data-efficient than dense per-point features. Key limitations are: noisy real-world scans challenge our manifold assumptions, Mahalanobis confidence degrades under distribution shift, and open-vocabulary generalization is limited to categories similar to the training data. Future work should extend this to articulated objects and integrate part-level alignments into foundation 3D models for manipulation and generation. By bridging dense geometry and structured language, ALIGN-Parts enables scalable, semantically rich 3D asset creation.
Appendix A Experiments and Analysis
@@
Given the challenges inherent in semantic 3D part segmentation, we find that no current published work is directly comparable to our method. To enable rigorous evaluation, we introduce our own strong baseline detailed in Section˜A.1. While we do include comparisons against class-agnostic 3D part segmentation methods in this manuscript, it is important to note that these do not constitute an entirely fair benchmark for our approach. Most prior methods have been trained using proprietary, closed-source Objaverse-scale datasets, with specific data details and part annotations rarely disclosed publicly.
@@
Despite these constraints, our method achieves competitive or superior performance relative to existing baselines. It is reasonable to expect that - if provided with similar data volume and computational resources - ALIGN-Parts would further extend its advantage on standard metrics and benchmarks. Our design choices thus not only democratize research in 3D part segmentation but also highlight the promise of reproducibility, accessibility, and efficiency for large-scale semantic understanding in open 3D datasets.
A.1. PartField+MPNet baseline
@@
A.2. Part-Retrieval Comparison with Find3D
Figure 13. Part Retrieval Comparison with Find3D. We demonstrate text-driven part retrieval on two representative objects (airplane and motorbike) from Objaverse-General. Given natural language part queries (e.g., “body”, “wing”, “gas tank”, “wheel”), ALIGN-Parts identifies and retrieves spatially coherent point groups corresponding to each part. Compared to Find3D (left), our method produces more semantically and spatially consistent part retrievals by leveraging the hierarchical point $\rightarrow$ partlet $\rightarrow$ part label decomposition. This design encourages the discovery of well-connected, semantically meaningful regions rather than fragmented point clusters. Ground truth part segmentations (right) show the target labels. ALIGN-Parts achieves results that closely align with ground truth, validating the effectiveness of our partlet-based formulation for open-vocabulary part localization and retrieval.
@@ A.2. Part-Retrieval Comparison with Find3D
We present qualitative comparisons with Find3D on two representative 3D objects from the airplane and motorbike object classes in the Objaverse-General benchmark (part of our closed-vocabulary evaluation set), released by the Find3D authors. As shown in Figure 13, ALIGN-Parts consistently retrieves more spatially coherent and semantically meaningful part groups, demonstrating the effectiveness of our partlet-based design for part localization and retrieval tasks.

### v2

@@ 5 Limitations and Future Work
Future work will focus on mitigating this dependency by exploring self-supervised or weakly supervised formulations and by incorporating stronger 3D priors, for example from generative models or skeletal/medial representations. Another important direction is to reduce the current reliance on frozen PartField features by enabling full end-to-end training of the geometric feature extractor, which was not pursued here primarily due to computational constraints rather than methodological ones. Despite these issues, our framework is immediately usable by parties with abundant compute and proprietary 3D assets, who can scale ALIGN-Parts to richer, closed-source datasets and drive progress towards truly large-scale 3D scene understanding at the part level.
6 Discussion
@@
ALIGN-Parts reframes semantic 3D part segmentation as a set alignment problem, where Partlets trained via optimal transport jointly learn geometry and semantics without part count supervision. Unlike Find3D’s per-point alignments or PartField’s brittle clustering, our end-to-end matching produces fast and coherent named parts directly (Fig.˜4). LLM-generated affordance descriptions ("door handle is grasped to open a door") are important for disambiguating fine-grained parts that confuse simple part names, and MPNet’s handling of long-form text outperforms CLIP/SigLIP (Fig.˜3). Despite using 10× fewer points than PartField, we achieve superior fine-part localization (e.g., a scissors screw, Fig.˜7), confirming that semantic part-level representations are more data-efficient than dense per-point features. Key limitations are: noisy real-world scans challenge our manifold assumptions, Mahalanobis confidence degrades under distribution shift, and open-vocabulary generalization is limited to categories similar to the training data. Future work should extend this to articulated objects and integrate part-level alignments into foundation 3D models for manipulation and generation. By bridging dense geometry and structured language, ALIGN-Parts enables scalable, semantically rich 3D asset creation.
Supplementary
Appendix A Architecture
A.1 Dense Feature Fusion Module
The raw geometric features $\mathbf{f}_{i}^{g}$ and appearance features $\mathbf{f}_{i}^{a}$ capture complementary information: geometry encodes shape characteristics, while appearance provides texture and visual cues. We fuse these modalities through a bi-directional cross-attention module that operates on a $k$-nearest neighbor (KNN) graph to maintain computational tractability.
Relative Positional Bias. To incorporate 3D spatial structure through Fourier-encoded relative positional biases, for each neighbor pair $(i,j)$ where $j\in\mathcal{N}_{i}$, we compute:
$\mathbf{d}_{ij}$ $=\mathbf{x}_{j}-\mathbf{x}_{i}\in\mathbb{R}^{3}$ (26)
$\mathbf{f}_{ij}$ $=\mathbf{d}_{ij}\odot\omega\in\mathbb{R}^{3\times F}$ (27)
$\mathbf{h}_{ij}$ $=[\sin(\mathbf{f}_{ij}),\cos(\mathbf{f}_{ij})]\in\mathbb{R}^{3\times 2F}$ (28)
$\mathbf{b}_{ij}$ $=\text{MLP}([\mathbf{d}_{ij},\text{flatten}(\mathbf{h}_{ij})])\in\mathbb{R}^{H}$ (29)
Bi-Directional Cross-Attention.
Geometric-to-Appearance Direction:
$\mathbf{Q}_{p}^{i}$ $=\mathbf{W}_{q}^{p}\mathbf{f}_{i}^{g}\in\mathbb{R}^{H\times d_{h}}$ (30)
$\mathbf{K}_{a}^{ij}$ $=\mathbf{W}_{k}^{a}\mathbf{f}_{j}^{a}\in\mathbb{R}^{H\times d_{h}},\quad\forall j\in\mathcal{N}_{i}$ (31)
$\mathbf{V}_{a}^{ij}$ $=\mathbf{W}_{v}^{a}\mathbf{f}_{j}^{a}\in\mathbb{R}^{H\times d_{h}},\quad\forall j\in\mathcal{N}_{i}$ (32)
$\alpha_{pa}^{i,h,j}$ $=\frac{\exp((\mathbf{Q}_{p}^{i}[h]\cdot\mathbf{K}_{a}^{ij}[h])/\sqrt{d_{h}}+\mathbf{b}_{ij}[h])}{\sum_{j^{\prime}\in\mathcal{N}_{i}}\exp((\mathbf{Q}_{p}^{i}[h]\cdot\mathbf{K}_{a}^{ij^{\prime}}[h])/\sqrt{d_{h}}+\mathbf{b}_{ij^{\prime}}[h])}$ (33)
$\mathbf{z}_{p}^{i,h}$ $=\sum_{j\in\mathcal{N}_{i}}\alpha_{pa}^{i,h,j}\mathbf{V}_{a}^{ij}[h]$ (34)
$\mathbf{r}_{p}^{i}$ $=\mathbf{W}_{pa}\text{concat}_{h}[\mathbf{z}_{p}^{i,h}]\in\mathbb{R}^{d_{g}}$ (35)
Appearance-to-Geometric Direction:
Gated Fusion.
$\mathbf{g}_{p}^{i}$ $=\sigma(\mathbf{W}_{g}^{p}[\mathbf{f}_{i}^{g};\mathbf{r}_{p}^{i}])\in\mathbb{R}^{d_{g}}$ (36)
$\mathbf{g}_{a}^{i}$ $=\sigma(\mathbf{W}_{g}^{a}[\mathbf{f}_{i}^{a};\mathbf{r}_{a}^{i}])\in\mathbb{R}^{d_{a}}$ (37)
$\tilde{\mathbf{f}}_{i}^{g}$ $=\text{LayerNorm}(\mathbf{f}_{i}^{g}+\mathbf{g}_{p}^{i}\odot\mathbf{r}_{p}^{i})$ (38)
$\tilde{\mathbf{f}}_{i}^{a}$ $=\text{LayerNorm}(\mathbf{f}_{i}^{a}+\mathbf{g}_{a}^{i}\odot\mathbf{r}_{a}^{i})$ (39)
Final Projection.
$\mathbf{h}_{i}=\mathbf{W}_{2}\text{GELU}(\mathbf{W}_{1}\text{LayerNorm}([\tilde{\mathbf{f}}_{i}^{g};\tilde{\mathbf{f}}_{i}^{a}]))$ (40)
Figure 11: Part Retrieval Comparison with Find3D. We demonstrate text-driven part retrieval on two representative objects (airplane and motorbike) from Objaverse-General. Given natural language part queries (e.g., “body”, “wing”, “gas tank”, “wheel”), ALIGN-Parts identifies and retrieves spatially coherent point groups corresponding to each part. Compared to Find3D (left), our method produces more semantically and spatially consistent part retrievals by leveraging the hierarchical point $\rightarrow$ partlet $\rightarrow$ part label decomposition. This design encourages the discovery of well-connected, semantically meaningful regions rather than fragmented point clusters. Ground truth part segmentations (right) show the target labels. ALIGN-Parts achieves results that closely align with ground truth, validating the effectiveness of our partlet-based formulation for open-vocabulary part localization and retrieval.
Appendix B Experiments and Analysis
@@
Given the challenges inherent in semantic 3D part segmentation, we find that no current published work is directly comparable to our method. To enable rigorous evaluation, we introduce our own strong baseline detailed in Sec.˜B.1. While we do include comparisons against class-agnostic 3D part segmentation methods in this manuscript, it is important to note that these do not constitute an entirely fair benchmark for our approach. Most prior methods have been trained using proprietary, closed-source Objaverse-scale datasets, with specific data details and part annotations rarely disclosed publicly.
@@
Despite these constraints, our method achieves competitive or superior performance relative to existing baselines. It is reasonable to expect that, if provided with similar data volumes and computational resources, ALIGN-Parts would further extend its advantage on standard metrics and benchmarks. Our design choices thus not only democratize research in 3D part segmentation but also highlight the promise of reproducibility, accessibility, and efficiency for large-scale semantic understanding in open 3D datasets.
B.1 PartField+MPNet baseline
@@
Figure 12: 3D Shape Co-Segmentation Analysis. Left: clustering-based co-segmentation. Prior methods such as PartField perform dependent clustering by first segmenting a source shape via feature clustering and then using the resulting cluster means to initialize K-means on a target shape, implicitly enforcing part correspondence; this can break when the target has a different part count or geometry, causing errors such as the red backrest region bleeding into the seat on the target chair. Using the same dependent co-segmentation strategy with our fused BiCo features (BiCo Feature Dependent Clustering) yields improved transfers on moderately similar targets (Target A), but performance degrades on more challenging targets with greater variation in part structure (Target B). As an alternative, we apply independent clustering to Target C, where the target is segmented with source initialization, and clusters are matched post hoc by comparing source and target cluster centers, which proves more reliable for difficult co-segmentation cases. Middle and right: feedforward ALIGN-Parts. In contrast to all clustering-based variants, the proposed feedforward ALIGN-Parts model (middle) directly predicts part segmentation and names, achieving robust results across shapes with differing part counts and topologies, and eliminating any dependence on source shapes or explicit co-segmentation.
B.2 Part-Retrieval Comparison with Find3D
@@
We present qualitative comparisons with Find3D on two representative 3D objects from the airplane and motorbike object classes in the Objaverse-General benchmark [26] (part of our closed-vocabulary evaluation set). As shown in Fig.˜11, ALIGN-Parts consistently retrieves more spatially coherent and semantically meaningful part groups, demonstrating the effectiveness of our partlet-based design for part localization and retrieval tasks.
B.3 3D Shape Co-Segmentation and Part Label Transfer
Fig.˜12 shows results and analysis of 3D Shape Co-Segmentation using ALIGN-Parts (and the BiCo features) as compared to PartField. A classical approach to 3D part segmentation operates in a co-segmentation setting, where multiple shapes from the same category are jointly analyzed to establish consistent part correspondence. Prior methods, including PartField, employ what we call dependent clustering: they first segment a source shape using feature clustering, then initialize k-means clustering on a target shape using the source cluster centroids, implicitly enforcing part correspondence. While this strategy can succeed on geometrically similar shapes, it proves fragile when target shapes exhibit different part counts or topologies. For example, in Fig.˜12, dependent clustering on a moderately similar target chair (Target A) produces reasonable results, but fails dramatically on targets with substantial part variation (Target B), causing geometric boundaries to blur (e.g., the backrest merging incorrectly with the seat).
In contrast to both clustering-based paradigms, our proposed ALIGN-Parts adopts a fully feedforward, discriminative approach that predicts part segmentation masks and semantic labels jointly, without requiring source shape initialization or explicit co-segmentation. This design eliminates brittleness to part count variation and geometric mismatches, enabling robust generalization across shapes with diverse part structures and semantics. As demonstrated in Fig.˜12 (middle and right panels), ALIGN-Parts consistently produces accurate, semantically grounded part segmentations regardless of target shape complexity.
Appendix C Applications
Named 3D part decompositions enable a range of downstream applications beyond segmentation benchmarks. ALIGN-Parts can serve as a scalable annotation and taxonomy-normalization engine for dataset construction, complementing part-annotated resources such as PartNet, 3DCoMPaT/3DCoMPaT++, and GAPartNet, which rely on inconsistent or category-specific taxonomies. In robotics and embodied AI, semantic part names aligned with affordances (e.g., handle, support, hinge) provide a natural interface for manipulation and task planning, and can act as structured perceptual inputs to vision-language-action (VLA) models that require interpretable, compositional representations [1]. More broadly, named parts offer a compact, semantically grounded state abstraction for world models, where objects and scenes are represented as compositions of interacting components rather than monolithic entities  [11]. This representation is particularly relevant for scene understanding and visual question answering (VQA), where prior work on perceptual taxonomies and compositional reasoning argues that structured intermediate representations improve robustness, generalization, and interpretability [20]. In 3D reconstruction [29] and understanding, part-aware representations can support structured reconstruction, correspondence, and pose estimation [16] by enforcing semantic consistency across views and instances. Finally, in 3D editing and content creation, named parts enable targeted edits and modular control (e.g., modifying or replacing specific components), complementing recent part-based generative models that discover structure but lack open-vocabulary semantic naming [4]. By producing complete, permutation-consistent named part decompositions in a single feed-forward pass, ALIGN-Parts provides a general-purpose semantic substrate for scene reasoning, VQA, embodied decision-making, and 3D content manipulation without requiring prompts, predefined vocabularies, or task-specific pipelines.